# Towards Goal-oriented Prompt Engineering for Large Language Models: A Survey

### v2

@@ Abstract
Large Language Models (LLMs) have shown prominent performance in various downstream tasks and prompt engineering plays a pivotal role in optimizing LLMs’ performance. This paper, not only as an overview of current prompt engineering methods, but also aims to highlight the limitation of designing prompts based on an anthropomorphic assumption that expects LLMs to think like humans. From our review of 36 representative studies, we demonstrate that a goal-oriented prompt formulation, which guides LLMs to follow established human logical thinking, significantly improves the performance of LLMs. Furthermore, We introduce a novel taxonomy that categorizes goal-oriented prompting methods into five interconnected stages and we demonstrate the broad applicability of our framework. With four future directions proposed, we hope to further emphasize the power and potential of goal-oriented prompt engineering in all fields.
@@ 1 Introduction
Large Language Models (LLMs) have garnered significant interest with their ability to assimilate vast amounts of information from large corpora of data. They have demonstrated proficiency in open-domain dialogue Deng et al. (2023); Yu et al. (2023), reasoning Wei et al. (2022); Yao et al. (2023); Besta et al. (2023), planning Song et al. (2023); Liu et al. (2023a), and many more tasks. The predominant approach to leveraging LLMs for downstream tasks involves crafting tailored text prompts to tap into LLMs’ potential Liu et al. (2023c). However, the effectiveness of LLMs is contingent upon the prompting strategy employed, leading to performance disparities Zhao et al. (2021). Such a phenomenon has given rise to the field of prompt engineering, which investigates the optimal formulation of prompts for specific tasks.
To prompt LLMs appropriately, it is important to first understand the deviation in the human way of thinking and LLMs’ way of processing information. Humans are taught by experience to approach complex objectives by breaking down the main goal into more manageable sub-goals, a strategy supported by goal-setting theory Austin and Vancouver (1996). In interacting with LLMs, which often display human-like conversational abilities, we might be tempted to attribute human-like thought processes to them, expecting them to know from the start that they should decompose complex problems into simpler tasks. This anthropomorphic assumption does not hold for LLMs and can lead to unsatisfactory outcomes Abercrombie et al. (2023). On the other hand, as shown in our literature review of 36 representative studies, if we design the prompts from a goal-oriented perspective that guides LLMs to mimic human thinking, LLMs’ performance improves significantly. In the meantime, aligning LLM’s behavior with human logic allows more human-interpretable answers from LLMs and results in more effective and reliable human-computer interaction.
@@
While several surveys have explored LLM prompting, none have examined it from a goal-oriented perspective or explicitly summarized the task-agnostic internal steps involved in designing and refining prompting procedures. Some surveys summarize studies that focus on specific tasks or problems. Qiao et al. (2022); Huang and Chang (2022) focus solely on reasoning tasks and Wang et al. (2024) on planning tasks. Huang et al. (2023) deals with the hallucination problem of LLMs. Other studies focus on generally applicable prompting but different sub-fields. Amatriain (2024); Sahoo et al. (2024) simply listed out recent techniques without systematic categorization. Liu et al. (2023c) focus on the prompting framework for facilitating the interaction between LLMs with external worlds. Liu et al. (2023b) focuses on methods for learnable continuous prompt, while our work studies discrete prompt. As many of the will-functioned LLMs are not open-sourced (e.g. GPT-3.5, GPT4), discrete prompts became the focus of recent studies on LLM prompting and therefore we choose to focus on discrete prompting.
In this paper, we propose the first goal-oriented taxonomy of prompting methods, which classifies goal-oriented prompting methods into goal decomposition, action selection, action execution, sub-goal result evaluation, and sub-goal selection, as shown in Fig. 1. We further classify the surveyed methods by their targeted task and analyze their respective performance to provide readers with a better estimate of the current progress of goal-oriented prompt engineering. Finally, we offer four promising future directions, including hierarchical decomposition, efficiency problems, the synergy of stages, and broader applications.
@@ 2.1 Decomposing Goal into Sub-goal Sequences
Iterative decomposition generates a sub-goal, gets sub-goal results, and repeats this process with the knowledge of the previous sub-goal and sub-goal results. Chain-of-thought prompting (CoT) Wei et al. (2022) can be considered the first work that follows iterative decomposition. With either “Let’s think step by step.” Kojima et al. (2022), or intermediate reasoning steps in in-context examples in the prompt, LLMs can imitate the decomposition process and solve the problem step by step. Here, the intermediate reasoning steps can be considered as sub-goals, and they are sequentially connected to form a sub-goal sequence that leads to a correct output. CoT implicitly follows iterative decomposition as LLMs generate tokens in an autoregressive way, with the next sub-goal decided based on previous content.
There are also works that explicitly ask LLMs to follow the iterative decomposition. DecomP Khot et al. (2022) and Successive Prompting Dua et al. (2022) are two contemporary works that repeatedly ask sub-questions to get background knowledge for question answering tasks. The background knowledge required to answer the question is sub-goal to be achieved. Different from CoT which may generate sub-questions sequentially in one output, these two works repeatedly prompt LLMs to ask follow-up questions. Self-ask Press et al. (2022) improves the efficiency of DecomP and Successive prompting by designing a template for follow-up questions. With that template, LLMs can generate all essential questions and their answers by prompting once. Besides, in this paper, the authors empirically show that LLMs are often wrong when asking them to answer a complex question even if they know the true answer of all needed sub-questions. This finding indicates the significance of decomposing complex goals into simple sub-goals for LLMs.
@@
In contrast to iterative decomposition, plan-then-execute decomposition methods generate the sub-goal sequence all at once, which means that the latter sub-goal will not be affected by the result of the former ones. For example, Least-to-most prompting Zhou et al. (2022) only prompts LLMs two times, one for generating a plan to decompose the goal into a sub-goal sequence, and the other one for executing the plan. Plan-and-solve prompting Wang et al. (2023b) further improves the efficiency of Least-to-most prompting by merging the generation of the plan and execution of the plan into one output. DEPS Wang et al. (2023d) and GITM Zhu et al. (2023) are decomposition methods designed for Minecraft, an open-world game in which an agent can craft different items and tools. In this game, obtaining base materials that are needed to craft a target item can be viewed as sub-goals. While DEPS generates a plan to obtain required objects in sequence solely based on LLMs, GITM leverages a pre-defined sub-goal tree to help LLMs locate prerequisites more precisely.
@@ 2.2 Action Selection for Attaining Sub-goals
With sub-goals defined, another important step is action selection, choosing the most effective and valid actions to reach desired outcomes. CoT is a naive way for action selection where it is completely decided by LLMs themselves. However, due to the hallucination problem, actions generated by LLMs are often invalid. In this section, we classify the advanced action selection methods into two classes, constrain-then-select and select-then-mapping.
@@
Constrain-then-select method predefines an action space and then has LLMs select the actual action among the space. As in MWP Zhang et al. (2023), to deal with math word problems, the authors first employ an operation prediction module to predict the needed calculation operation (e.g. multiplication, summation). LLMs are then used to select the appropriate operands and complete the calculation. PEARL Sun et al. (2023) utilizes constrained action space for question answering over long documents. The authors predefine actions like “Finding the definition of A”, “Compare A and B”, and “Summarize A”. DecomP Khot et al. (2022) selects actions from a set like “split” and “merge” for k-th letter concatenation, a symbolic reasoning task. SayPlan Rana et al. (2023) is for robot planning. Given a task instruction, the authors apply semantic search to first identify a task-relevant subgraph from the whole 3D scene graph, then have LLMs plan based on the subgraph. In dialogue systems, RLP Fischer (2023) allows the user to set a mental state of LLM, such as beliefs or desires, to instruct action selection. Such a setting improves the context richness, coherence, and interactivity of LLM-based conversational systems. Similar to PEARL, ProCoT Deng et al. (2023) forms pre-defined action sets covering query clarification, topic transition, and negotiation strategy for dialogue systems.
@@
Different from constrain-then-select, select-then-mapping first uses LLMs to generate actions solely based on their knowledge and then maps the generated action to the most similar one in the valid action space. Zero-shot planners Huang et al. (2022a), SALP Gramopadhye and Szafir (2023), and Re-Prompting Raman et al. (2022) aim to solve agent planning problems in an interactive environment. In such virtual environments, there are only a few admissible actions that can be applied. However, the actions produced naively by LLMs often cannot be mapped exactly to those executable actions. To remedy this, researchers employ a text-embedding language model to translate LLM-generated actions into the most similar admissible actions by calculating cosine similarity.
@@ 2.3 Executing Actions to Get Sub-goal Results
In CoT, LLMs rely on their knowledge for action execution to get the sub-goal result. Due to the hallucination problem, the results are sometimes wrong. For example, even if LLMs correctly select multiplication as the action to solve a math problem, the calculation result can be wrong. Besides, as Khot et al. (2022) claimed, some tasks may not be feasible to solve using only one LLM. With such consideration, some works leverage external tools to guarantee a precise sub-goal result.
Some studies specifically study one task Khot et al. (2022); Chen et al. (2022); Wang et al. (2023c). To answer open-domain questions, DecomP Khot et al. (2022) applies an ElasticSearch-based retrieval system to retrieve knowledge from certain knowledge bases like Wikipedia. Program-of-Thoughts (PoT) Chen et al. (2022) first translates reasoning processes into Python codes and then runs them in an interpreter. Recmind Wang et al. (2023c) uses SQL to search for the history of a user in the database in recommender systems.
There are additional studies that aim for multi-tasking. Toolformer Schick et al. (2023) specifically studied which APIs to call and when to call them when using external tools. In their paper, several tools are adopted, including a calculator, a calendar, a BM25-based Wikipedia search engine, and a fine-tuned language model (LM). The authors pointed out that small LMs finetuned for specific tasks can outperform some LLMs on certain task. Following the idea of Toolformer, HuggingGPT Shen et al. (2023) further extends the available LMs to the scale of all LMs on open-sourced machine learning communities like HuggingFace. Specifically, in their paper, they evaluate HuggingGPT on nine types of NLP tasks, nine types of CV tasks, four types of audio tasks, and two types of video tasks. Additionally, with the development of open-sourced communities, this work demonstrates the potential of achieving all the sub-goals that can be described in language by leveraging external tools.
@@ 2.4 Evaluating Sub-goal Results
LLM Self-refine Madaan et al. (2023)
SelfCheck Miao et al. (2023)
Reflexion Shinn et al. (2023)
Environment SayPlan Rana et al. (2023)
Re-prompt Raman et al. (2022)
Inner Monologue Huang et al. (2022b)
GITM  Zhu et al. (2023)
DEPS  Wang et al. (2023d)
VLM DEPS  Wang et al. (2023d)
LLM-Planner Song et al. (2023)
Inner Monologue Huang et al. (2022b)
Program Executor Self-debug Chen et al. (2023)
INTERVENOR Wang et al. (2023a)
@@
Human Inner Monologue Huang et al. (2022b)
DEPS  Wang et al. (2023d)
@@
In CoT, there is no evaluation process for the sub-goal results. The results generated by LLMs are assumed to be correct by default. Answering by LLMs, sub-goal results also suffer from the hallucination problem. In addition, errors may propagate along the pipeline and result in huge deviations from the correct answer. Thus, it is necessary to evaluate the sub-goal results at each step and correct them in time. Note that the methods we review in this section can be applied to evaluate the overall goal achievement as well. The types of feedback source and their related works are summarized in Table 1.
@@
Evaluate with LLMs.
@@
Self-refine Madaan et al. (2023) is the first work addressing the evaluation of sub-goal results. After getting the output of LLMs, Self-refine feeds the output combined with task-specific prompts to the same LLM to get feedback for the output. Then, the feedback and initial output are fed together to LLMs for output refinement. The refinement process iterates with all of the past feedback and refinement appended to the prompt. For the correction process of the following works in this section, they all follow Self-refine to correct sub-goal results, if not specified. SelfCheck Miao et al. (2023) decomposes the Self-refine evaluation into finer-grained processes. First, actions and sub-goal results are put to LLMs to summarize their intention. Then, the goal combined with all previous sub-goals, selected actions, and sub-goal results serves as another input to let LLMs extract useful information from them. Finally, only being provided with useful information and summarized intention, LLMs generate the sub-goal result again. This result will be compared with the original result. If the regeneration result supports the original sub-goal result, the original result is considered correct, and vice versa. In Reflexion Shinn et al. (2023), the authors try using another instantiation of an LLM to obtain a score for the sub-goal result.
Evaluate with external evaluators.
@@
In contrast to self-evaluation, we could seek the help of external evaluators for providing feedback, which is similar to but different from the idea of applying external tools for action execution.
@@
Some methods Chen et al. (2023); Wang et al. (2023a); Rana et al. (2023); Raman et al. (2022) leverage the error messages that are pre-built into either the programming languages or the program of virtual environments to evaluate sub-goal results. Self-debug Chen et al. (2023) and INTERVENOR Wang et al. (2023a) leverages unit tests and executors to gain feedback on generated code snippets. The authors also empirically show that leveraging unit test error messages leads to superior performance than LLM’s self-reflection. This finding indicates that external evaluators may be more beneficial than self-reflection thanks to the more precise feedback. SayPlan Rana et al. (2023) and Re-prompting Raman et al. (2022) directly leverage the signs of success or failure with error messages from the environment as feedback to revise their actions. GITM Zhu et al. (2023), a method proposed for Minecraft, designs a structured feedback template to get detailed knowledge of the current situation. All of the information can be obtained from Minecraft API returns.
@@
The quality of error messages (e.g. detail) is influential to the evaluating process, but sometimes it is hard to define error messages in advance. To further address it, LLM-Planner Song et al. (2023) explores the possibility of using Vision-Language Models (VLMs) to provide feedback by describing the environment after confirming that the action has failed. Inner Monologue Huang et al. (2022b) is a combination of SayPlan and LLM-Planner, where VLMs additionally perform the task of visual question answering, which enables a finer-grained human-computer interaction to get valuable information. Similar to Inner Monologue, DEPS Wang et al. (2023d), designed for Minecraft, also gets feedback from both VLMs and environments. DEPS further introduces a step to ask LLMs to explain the reason according to the feedback. Reflexion Shinn et al. (2023) proposes various evaluators to handle different tasks. It has reward functions for reasoning tasks based on exact match grading and heuristic functions tailored to specific criteria for decision-making tasks. We refer to them as heuristic rules. Note that the correction process of SayPlan, Re-prompting, LLM-Planner, Inner Monologue, DEPS, and GITM is different from Self-refine. Instead of executing the action to get the sub-goal result again, these methods restart from action selection.

### v3

@@ Abstract
Large Language Models (LLMs) have shown prominent performance in various downstream tasks and prompt engineering plays a pivotal role in optimizing LLMs’ performance. This paper, not only as an overview of current prompt engineering methods, but also aims to highlight the limitation of designing prompts based on an anthropomorphic assumption that expects LLMs to think like humans. From our review of 50 representative studies, we demonstrate that a goal-oriented prompt formulation, which guides LLMs to follow established human logical thinking, significantly improves the performance of LLMs. Furthermore, We introduce a novel taxonomy that categorizes goal-oriented prompting methods into five interconnected stages and we demonstrate the broad applicability of our framework. With four future directions proposed, we hope to further emphasize the power and potential of goal-oriented prompt engineering in all fields.
@@ 1 Introduction
Large Language Models (LLMs) have garnered significant interest with their ability to assimilate vast amounts of information from large corpora of data. They have demonstrated proficiency in open-domain dialogue Deng et al. (2023); Yu et al. (2023a), reasoning Wei et al. (2022); Yao et al. (2023b); Besta et al. (2023), planning Song et al. (2023); Liu et al. (2023a), and many more tasks. The predominant approach to leveraging LLMs for downstream tasks involves crafting tailored text prompts to tap into LLMs’ potential Liu et al. (2023d). However, the effectiveness of LLMs is contingent upon the prompting strategy employed, leading to performance disparities Zhao et al. (2021). Such a phenomenon has given rise to the field of prompt engineering, which investigates the optimal formulation of prompts for specific tasks.
To prompt LLMs appropriately, it is important to first understand the deviation in the human way of thinking and LLMs’ way of processing information. Humans are taught by experience to approach complex objectives by breaking down the main goal into more manageable sub-goals, a strategy supported by goal-setting theory Austin and Vancouver (1996). In interacting with LLMs, which often display human-like conversational abilities, we might be tempted to attribute human-like thought processes to them, expecting them to know from the start that they should decompose complex problems into simpler tasks. This anthropomorphic assumption does not hold for LLMs and can lead to unsatisfactory outcomes Abercrombie et al. (2023). On the other hand, as shown in our literature review of 50 representative studies, if we design the prompts from a goal-oriented perspective that guides LLMs to mimic human thinking, LLMs’ performance improves significantly. In the meantime, aligning LLM’s behavior with human logic allows more human-interpretable answers from LLMs and results in more effective and reliable human-computer interaction.
@@
While several surveys have explored LLM prompting, none have examined it from a goal-oriented perspective or explicitly summarized the task-agnostic internal steps involved in designing and refining prompting procedures [^1] We discuss related surveys in Appendix A.. In this paper, we propose the first goal-oriented taxonomy of prompting methods, which classifies goal-oriented prompting methods into goal decomposition, action selection, action execution, sub-goal result evaluation, and sub-goal selection, as shown in Fig. 1. We further classify the surveyed methods by their targeted task and analyze their respective performance to provide readers with a better estimate of the current progress of goal-oriented prompt engineering. Finally, we offer four promising future directions, including hierarchical decomposition, efficiency problems, the synergy of stages, and broader applications.
2 Preliminary
In this section, we provide a definition of goal-oriented prompt engineering. Before define goal-oriented prompting, we can first define standard prompting, the process of prompting LLM $Pr_{LLM}$ for a prediction on a given goal as:
$Pr(\mathcal{P}|\mathcal{G})=\prod_{i=1}^{|\mathcal{P}|}Pr_{LLM}(p_{i}|\mathcal{G},p_{<i})$ (1)
By introducing additional variables sub-goal $g$, action $a$, and sub-goal result $r$, we formally define goal-oriented prompt engineering as the application of a multi-stage strategy to guide LLMs in problem-solving by adhering to established human logic, as shown in Fig. 1.
In goal-oriented prompt engineering, the overall goal $\mathcal{G}$ is first decomposed into sub-goals $\{g_{1},g_{2},...,g_{n}\}$ (Fig.1(1)), then actions $\{a_{1},a_{2},...,a_{n}\}$ are selected to achieve corresponding sub-goals (Fig.1(2)). The final prediction is based on valuable sub-goal results $\{r_{1},r_{2},...,r_{n}\}$ (Fig.1(5)). The process could be described as:
$Pr(\mathcal{P}|\mathcal{G})$ $=\sum_{i=1}^{n}Pr(P|r_{i},a_{i},g_{i},\mathcal{G})$ (2)
$Pr(r_{i}|a_{i},g_{i},\mathcal{G})Pr(a_{i}|g_{i},\mathcal{G})Pr(g_{i}|\mathcal{G})$
If LLMs are applied for the implementation of a stage, we could follow Eq. 1 to get the output. We additionally introduce the action execution (Fig.1(3)) and sub-goal evaluation stages (Fig.1(4)) to ensure the correctness of sub-goal results.
@@ 3.1 Decomposing Goal into Sub-goal Sequences
Iterative decomposition generates a sub-goal, gets sub-goal results, and repeats this process with the knowledge of the previous sub-goal and sub-goal results. Chain-of-thought prompting (CoT) Wei et al. (2022) can be considered the first work that follows iterative decomposition. With either “Let’s think step by step.” Kojima et al. (2022) in the prompt, or intermediate reasoning steps in in-context examples, LLMs can imitate the decomposition process and solve the problem step by step. Here, the intermediate reasoning steps can be considered as sub-goals, and they are sequentially connected to form a sub-goal sequence. CoT implicitly follows iterative decomposition as LLMs generate tokens in an autoregressive way, with the next sub-goal decided based on previous content.
There are also works that explicitly ask LLMs to follow the iterative decomposition. DecomP Khot et al. (2022), Successive Prompting Dua et al. (2022), and Recursion-of-Thought (RoT) Lee and Kim (2023) are three works that repeatedly prompt LLMs to ask follow-up sub-questions for background knowledge. Self-ask Press et al. (2022a) improves efficiency by designing a template for follow-up questions in advance. With that template, LLMs can generate all essential questions and their answers by prompting once. Besides, in this paper, the authors empirically show that LLMs are often wrong when asking them to answer a complex question even if they know the true answer of all needed sub-questions. This finding indicates the significance of decomposing complex goals into simple sub-goals for LLMs.
@@
In contrast to iterative decomposition, plan-then-execute decomposition methods generate the sub-goal sequence all at once, which means that the latter sub-goal will not be affected by the result of the former ones. For example, Least-to-most prompting Zhou et al. (2022) only prompts LLMs two times, one for generating a plan to decompose the goal into a sub-goal sequence, and the other one for executing the plan. Plan-and-solve prompting Wang et al. (2023d) further improves the efficiency of Least-to-most prompting by merging the generation of the plan and execution of the plan into one output. DEPS Wang et al. (2023f) and GITM Zhu et al. (2023) are decomposition methods designed for Minecraft, an open-world game in which an agent can craft different items and tools. In this game, obtaining base materials that are needed to craft a target item can be viewed as sub-goals. While DEPS generates a plan to obtain required objects in sequence solely based on LLMs, GITM leverages a pre-defined sub-goal tree to help LLMs locate prerequisites more precisely.
@@ 3.2 Action Selection for Attaining Sub-goals
With sub-goals defined, another important step is action selection, choosing effective and valid actions to reach desired outcomes. CoT is a naive way for action selection where it is completely decided by LLMs themselves. However, due to the hallucination problem, actions from LLMs are often invalid. In this section, we classify the advanced action selection methods into two classes, constrain-then-select and select-then-mapping.
@@
Constrain-then-select method predefines an action space and then has LLMs select the actual action among the space. As in MWP Zhang et al. (2023), to deal with math word problems, the authors first employ an operation prediction module to predict the needed calculation operation (e.g. multiplication, summation). LLMs are then used to select the appropriate operands and complete the calculation. PEARL Sun et al. (2023) utilizes constrained action space like “Finding the definition of A”, “Compare A and B”, and “Summarize A” for question answering over long documents. DecomP Khot et al. (2022) selects actions from a set like “split” and “merge” for k-th letter concatenation, a symbolic reasoning task. In robot planning, given a task instruction, SayPlan Rana et al. (2023) applies semantic search to first identify a task-relevant subgraph from the whole 3D scene graph, which makes it easier for LLMs to plan based on the subgraph. In dialogue systems, RLP Fischer (2023), SAFARI Wang et al. (2023b), and Cue-CoT Wang et al. (2023c) set the mental states of the LLM and the user, respectively, such as beliefs or desires, to guide action selection. Such a setting improves the context richness, coherence, and interactivity of LLM-based conversational systems. Similar to PEARL, ProCoT Deng et al. (2023) forms pre-defined action sets covering query clarification, topic transition, and negotiation strategy for dialogue systems.
@@
Different from constrain-then-select, select-then-mapping first uses LLMs to generate actions solely based on their knowledge and then maps the generated action to the most similar one in the valid action space. Zero-shot planners Huang et al. (2022b), SALP Gramopadhye and Szafir (2023), and Re-Prompting Raman et al. (2022) aim to solve agent planning problems in an interactive environment. In such virtual environments, there are only a few admissible actions that can be applied. However, the actions produced naively by LLMs often cannot be mapped exactly to those executable actions. To remedy this, researchers employ a text-embedding language model to translate LLM-generated actions into the most similar admissible actions by calculating cosine similarity.
@@ 3.3 Executing Actions to Get Sub-goal Results
In CoT, LLMs rely on their knowledge for action execution to get the sub-goal result. Limited by hallucination, the results are sometimes wrong. For example, even if LLMs correctly select multiplication as the action to solve a math problem, the calculation result can be wrong. As Khot et al. (2022) claimed, some tasks may not be feasible to solve using only LLMs. With such consideration, some works leverage external tools to guarantee a precise sub-goal result.
Some studies specifically study one task Khot et al. (2022); Chen et al. (2022); Wang et al. (2023e). To answer open-domain questions, DecomP Khot et al. (2022) applies an ElasticSearch-based retrieval system to retrieve knowledge from certain knowledge bases like Wikipedia. Program-of-Thoughts (PoT) Chen et al. (2022), Faithful CoT Lyu et al. (2023), PAL Gao et al. (2023), MathPrompter Imani et al. (2023), LINC Olausson et al. (2023), and Logic-LM Pan et al. (2023) translate reasoning processes into executable codes and then run them in an interpreter. In recommender systems, Recmind Wang et al. (2023e) and InteRecAgent Huang et al. (2023b) use SQL to search for the history of a user in the database while DOKE Yao et al. (2023a) achieves it through knowledge graphs.
There are additional studies that aim for multi-tasking. Toolformer Schick et al. (2023) specifically studied which APIs to call and when to call them when using external tools. In their paper, several tools are adopted, including a calculator, a calendar, a BM25-based Wikipedia search engine, and a fine-tuned language model (LM). The authors pointed out that small LMs finetuned for specific tasks can outperform some LLMs on certain task. Following the idea of Toolformer, HuggingGPT Shen et al. (2023) further extends the available LMs to the scale of all LMs on open-sourced machine learning communities like HuggingFace. Specifically, in their paper, they evaluate HuggingGPT on nine types of NLP tasks, nine types of CV tasks, four types of audio tasks, and two types of video tasks. With the development of open-sourced communities, this work can potentially achieve all the sub-goals that can be described in language.
@@ 3.4 Evaluating Sub-goal Results
LLM Self-refine Madaan et al. (2023), SelfCheck Miao et al. (2023), Reflexion Shinn et al. (2023), REFINER Paul et al. (2024)
Environment SayPlan Rana et al. (2023), Re-prompt Raman et al. (2022), Inner Monologue Huang et al. (2022c), GITM Zhu et al. (2023), DEPS Wang et al. (2023f)
VLM DEPS  Wang et al. (2023f), LLM-Planner Song et al. (2023), Inner Monologue Huang et al. (2022c)
Interpreter Self-debug Chen et al. (2023), INTERVENOR Wang et al. (2023a), CRITIC Gou et al. (2023), MAF Nathani et al. (2023)
@@
Human Inner Monologue Huang et al. (2022c), DEPS Wang et al. (2023f)
Search Engine CRITIC Gou et al. (2023), Verify-and-Edit Zhao et al. (2023)
@@
In CoT, the results generated by LLMs are assumed to be correct hence there is no evaluation process for the sub-goal results. The negative impact of hallucination becomes more severe for sub-goal results because errors may propagate along the pipeline and result in huge deviations from the correct answer. Thus, it is necessary to get feedback for sub-goal results at each step and correct them in time through resampling. The types of feedback sources are summarized in Table 1. Note that the methods we review in this section can be applied to evaluate the overall goal achievement as well.
@@
Feedback from LLMs.
@@
Self-refine Madaan et al. (2023); Huang et al. (2022a) is the first work addressing the evaluation of sub-goal results. After getting the output of LLMs, Self-refine feeds the output combined with task-specific prompts to the same LLM to get feedback for the output. Then, the feedback and initial output are fed together to LLMs for output refinement. The refinement process iterates with all of the past feedback and refinement appended to the prompt. SelfCheck Miao et al. (2023) decomposes the Self-refine evaluation into finer-grained processes. First, actions and sub-goal results are put to LLMs to summarize their intention. Then, the goal combined with all previous sub-goals, selected actions, and sub-goal results serves as another input to let LLMs extract useful information from them. Finally, only being provided with useful information and summarized intention, LLMs generate the sub-goal result again. This result will be compared with the original result. If the regeneration result supports the original sub-goal result, the original result is considered correct, and vice versa. In Reflexion Shinn et al. (2023) and REFINER Paul et al. (2024), the authors try using another instantiation of an LLM to obtain feedback for the sub-goal result.
Feedback from external tools.
@@
In contrast to self-evaluation, we could seek the help of external evaluators to provide feedback.
@@
Some methods Chen et al. (2023); Wang et al. (2023a); Rana et al. (2023); Raman et al. (2022) leverage the error messages that are pre-built into either the programming languages or the program of virtual environments to evaluate sub-goal results. Self-debug Chen et al. (2023), MAF Nathani et al. (2023) and INTERVENOR Wang et al. (2023a) leverages unit tests and executors to gain feedback on generated code snippets. The authors also empirically show that leveraging unit test error messages leads to superior performance than LLM’s self-reflection. This finding indicates that external evaluators may be more beneficial than self-reflection thanks to the more precise feedback Gou et al. (2023). SayPlan Rana et al. (2023) and Re-prompting Raman et al. (2022) directly leverage the signs of success or failure with error messages from the environment as feedback to revise their actions. GITM Zhu et al. (2023) designs a structured feedback template to get detailed knowledge of the current situation and all of the information can be obtained from environment API returns.
@@
However, sometimes it is hard to define error messages in advance. To further address it, LLM-Planner Song et al. (2023) explores the possibility of using Vision-Language Models (VLMs) to provide feedback by describing the environment. Inner Monologue Huang et al. (2022c) and DEPS Wang et al. (2023f) are combinations of SayPlan and LLM-Planner, where VLMs additionally perform visual question answering to enable a finer-grained human-computer interaction for valuable information. Reflexion Shinn et al. (2023) proposes various heuristic functions to score different tasks (e.g. exact match for reasoning). CRITIC Gou et al. (2023) and Verify-and-Edit Zhao et al. (2023) employ search engines to obtain true information as feedback from the web.